# Enhancing Safety of Foundation Models for Visual Navigation through Collision Avoidance via Repulsive Estimation

### v2

@@ Abstract
We propose CARE (Collision Avoidance via Repulsive Estimation), a plug-and-play module that enhances the safety of vision-based navigation without requiring additional range sensors or fine-tuning of pretrained models. While recent foundation models using only RGB inputs have shown strong performance, they often fail to generalize in out-of-distribution (OOD) environments with unseen objects or variations in camera parameters (e.g., field of view, pose, or focal length). Without fine-tuning, these models may generate unsafe trajectories that lead to collisions, requiring costly data collection and retraining. CARE addresses this limitation by seamlessly integrating with any RGB-based navigation system that outputs local trajectories, dynamically adjusting them using repulsive force vectors derived from monocular depth maps. We evaluate CARE by combining it with state-of-the-art vision-based navigation models across multiple robot platforms. CARE consistently reduces collision rates (up to 100%) without sacrificing goal-reaching performance and improves collision-free travel distance by up to 10.7× in exploration tasks.
@@ 2 Related Work
VLM-based navigation approaches such as LM-Nav [5], LFG [15], and NavGPT [3] leverage pretrained VLMs to interpret natural language commands and perform high-level planning without task-specific supervision. VLMaps [4] integrates vision-language features into 3D spatial maps, while Obstructed VLN [16] and knowledge-enhanced navigation [17] aim to improve robustness to domain shifts. However, these methods mainly focus on high-level semantic reasoning and goal specification, assuming idealized low-level controllers. As a result, safety-critical aspects such as collision avoidance remain largely unaddressed.
@@ 3 Collision Avoidance via Repulsive Estimation (CARE)
0:  RGB observation $I$, trajectory $\mathcal{T}$, depth estimation model $\mathcal{D}$
@@
1:  $\mathcal{O}\leftarrow\textit{ConstructTopDownObstacleMap}(I,\mathcal{D})$ // Stage 1: Top-down range estimation
2:  $\theta_{\text{rep}}\leftarrow\textit{EstimateRepulsiveDirection}(\mathcal{T},\mathcal{O})$ // Stage 2: Repulsive force estimation
3:  $\mathcal{T}^{\prime}\leftarrow\textit{RotateTrajectory}(\mathcal{T},\theta_{\text{rep}})$
4:  $\theta_{\text{des}}\leftarrow\textit{ComputeDesiredHeading}(\mathcal{T}^{\prime})$ // Stage 3: Safety-enhancing mechanism
@@ 3.1 Top-Down Range Estimation
CARE first constructs a top-down local map (see Figure 3(a)) by predicting the scene geometry from a depth image predicted from an RGB observation through the function ConstructTopDownObstacleMap$(I,\mathcal{D})$ (see Algorithm 1, Line 1).
@@ 3.2 Repulsive Force Estimation and Trajectory Adjustment
The repulsive force at each waypoint is computed via EstimateRepulsiveDirection$(\mathcal{T},\mathcal{O})$ (Algorithm 1, Line 2), following the same formulation used in a recent work on APF with limited sensing range [26]:
@@
$k^{\star}=\arg\max_{k\in\{1,\ldots,K\}}\|\mathbf{F}_{\text{rep}}(\mathbf{p}_{k})\|_{2}.$ (2)
@@
Finally, RotateTrajectory$(\mathcal{T},\theta_{\text{rot}})$ (Algorithm 1, Line 3) applies a 2D rotation matrix $R\in\mathbb{R}^{2\times 2}$ to each waypoint, yielding the adjusted trajectory $\mathcal{T}^{\prime}=\{\mathbf{p}^{\prime}_{1},\ldots,\mathbf{p}^{\prime}_{K}\}$, where each waypoint is updated as $\mathbf{p}^{\prime}_{k}=R(\theta_{\text{rot}})\cdot\mathbf{p}_{k}$. This rotation steers the trajectory away from nearby obstacles while preserving its overall direction toward the goal (see Figure 3(b)).
@@ 4 Experiments
We comprehensively evaluate CARE across unseen real environments and different robot platforms to validate our primary claim: CARE enhances collision avoidance capabilities of pretrained vision-based navigation models without fine-tuning, while preserving original navigation performance. To validate the generalizability of CARE, we deploy our method on three mobile robot platforms with different camera configurations: (i) LoCoBot with a 170°170°170  ° FOV fisheye camera [^1] The camera setup is consistent with that used in the baseline models NoMaD and ViNT., (ii) Clearpath TurtleBot4 equipped with an 89.5°89.5°89.5  ° FOV OAK-D Pro camera, and (iii) DJI RoboMaster S1 with a 120°120°120  ° FOV camera. All robots operate under consistent velocity constraints (maximum linear velocity $v_{\text{max}}$ of $0.2\,\mathrm{m/s}$ and angular velocity $w_{\text{max}}$ of $0.8\,\mathrm{rad/s}$) for comparable evaluation. We optimize CARE parameters experimentally for each platform while maintaining consistent evaluation across all platforms. Detailed robot-specific parameters are provided in the Appendix.
@@ 4.1 Undirected Exploration task in Unseen Environments
Figure 5 and Table 1 present our exploration experiment results. Across all settings, CARE enables robots to travel 2.9×2.9$\times$2.9 × to 10.7×10.7$\times$10.7 × longer before the first collision compared to NoMaD and ViNT without CARE. These performance variations correlated with the camera specifications of each configuration and the resulting depth estimation quality.
@@
LoCoBot with CARE exhibits both the longest distance traveled and the most significant improvement (10.7×10.7$\times$10.7 ×). This performance increase can be attributed to two main factors: the wide-angle fisheye camera enables extensive environmental perception, effectively detecting peripheral obstacles; and NoMaD, which has been predominantly trained on fisheye camera data, performs better with the camera setup of LoCoBot, which CARE further enhances.
TurtleBot4 with CARE exhibits moderate improvement ($\times3$) despite having the lowest distance traveled. This enhancement is partly due to its camera mounting position. When navigating corners, the rear-mounted camera maintains visibility of obstacles throughout the turn, allowing CARE to make safer adjustments. However, its narrow FOV still limits overall performance, particularly when approaching walls head-on.
RoboMaster without CARE achieves higher distance traveled due to its wider FOV, but shows the smallest improvement (2.9×2.9$\times$2.9 ×) with CARE. This limited gain stems from its front-mounted camera. During turning, the camera moves past obstacles before the body of the robot completes the turn, creating blind spots where side collisions can occur without detection. While the wider FOV achieves relatively higher distance traveled, it cannot fully compensate for these cornering blind spots.
@@ 4.2 Image Goal-Conditioned Navigation with Unseen Obstacles
We evaluate NoMaD and ViNT and their CARE-integrated variants in image goal-conditioned navigation scenarios involving unseen obstacles with the three robot platforms. We first establish a topological graph in a corridor environment to enable global planning. During testing, we place several OOD obstacles not represented in the topological map, creating scenarios that would typically require model fine-tuning to handle appropriately. Table 2 presents comprehensive results comparing the baseline navigation models (NoMaD and ViNT) against the same models integrated with CARE. All experiments are conducted across 10 distinct obstacle settings, with randomized positions for each trial. [^3] Detailed information about obstacle placements for each trial is available in the appendix.
@@
Navigation with dynamic obstacles. We conducted additional tests in the same environment as the goal-conditioned experiments (see Figure 4(b)), where up to three people abruptly crossed the robot path during navigation. A trial was considered successful if the robot reached the goal without collision. Unlike prior works, we introduced dynamic obstacles that appeared suddenly from outside the camera field of view. NoMaD often failed to react, while CARE-integrated robots consistently exhibited reactive stopping and safe avoidance, demonstrating the ability to handle unpredictable dynamic obstacles where vision-only models without fine-tuning often fail to generalize.
These results demonstrate the consistent ability of CARE to enhance navigation safety across various robot platforms without additional training. While performance varies based on camera configuration, CARE considerably improves collision avoidance in all tested scenarios, augmenting the capabilities of baseline models to handle complex environments with OOD objects. Supplementary information of these experiments are provided in the Appendix A.1.
@@ B.3 Common Parameters for CARE
$\epsilon$ (m) $-0.05$ Vertical offset for filtering out ceiling points during top-down projection. A negative value indicates upward exclusion

### v4

# CARE: Enhancing Safety of Visual Navigation through Collision Avoidance via Repulsive Estimation

@@ Abstract
We propose CARE (Collision Avoidance via Repulsive Estimation) to improve the robustness of learning-based visual navigation methods. Recently, visual navigation models, particularly foundation models, have demonstrated promising performance by generating viable trajectories using only RGB images. However, these policies can generalize poorly to environments containing out-of-distribution (OOD) scenes characterized by unseen objects or different camera setups (e.g., variations in field of view, camera pose, or focal length). Without fine-tuning, such models could produce trajectories that lead to collisions, necessitating substantial efforts in data collection and additional training. To address this limitation, we introduce CARE, an attachable module that enhances the safety of visual navigation without requiring additional range sensors or fine-tuning of pretrained models. CARE can be integrated seamlessly into any RGB-based navigation model that generates local robot trajectories. It dynamically adjusts trajectories produced by a pretrained model using repulsive force vectors computed from depth images estimated directly from RGB inputs. We evaluate CARE by integrating it with state-of-the-art visual navigation models across diverse robot platforms. Real-world experiments show that CARE significantly reduces collisions (up to 100%) without compromising navigation performance in goal-conditioned navigation, and further improves collision-free travel distance (up to 10.7×) in exploration tasks. Project page: https://airlab-sogang.github.io/CARE/
@@ 2 Related Work
VLM-based navigation approaches such as LM-Nav [5], LFG [15], and NavGPT [3] leverage pretrained VLMs to interpret natural language commands and perform high-level planning without task-specific supervision. VLMaps [4] integrates vision-language features into 3D spatial maps, while Obstructed VLN [16] and Knowledge-Enhanced Scene Understanding [17] aim to improve robustness to domain shifts. However, these methods mainly focus on high-level semantic reasoning and rely on discrete or graph-based action spaces, leaving safety-critical aspects such as collision avoidance largely unaddressed.
@@ 3 Collision Avoidance via Repulsive Estimation (CARE)
0: RGB observation II, trajectory $\mathcal{T}$, depth estimation model $\mathcal{D}$
@@
[…]
@@ 3.1 Top-Down Range Estimation
[…]
@@ 3.2 Repulsive Force Estimation and Trajectory Adjustment
[…]
@@
$k^{\star}=\mathrm{argmax}_{k\in\{1,\ldots,K\}}\|\mathbf{F}_{\text{rep}}(\mathbf{p}_{k},\mathcal{O})\|_{2}.$ (2)
@@
[…]
@@ 4 Experiments
We comprehensively evaluate CARE across unseen real environments and different robot platforms to validate our primary claim: CARE enhances collision avoidance capabilities of pretrained vision-based navigation models without fine-tuning, while preserving original navigation performance. To validate the generalizability of CARE, we deploy our method on three mobile robot platforms with different camera configurations: (i) LoCoBot with a 170° FOV fisheye camera [^1] The camera setup is consistent with that used in the baseline models NoMaD and ViNT., (ii) Clearpath TurtleBot4 equipped with an 89.5° FOV OAK-D Pro camera, and (iii) DJI RoboMaster S1 with a 120° FOV camera. All robots operate under consistent velocity constraints (maximum linear velocity $v_{\text{max}}$ of $0.2\,\mathrm{m/s}$ and angular velocity $w_{\text{max}}$ of $0.8\,\mathrm{rad/s}$) for comparable evaluation. We optimize CARE parameters experimentally for each platform while maintaining consistent evaluation across all platforms. Detailed robot-specific parameters are provided in the Appendix.
@@ 4.1 Undirected Exploration task in Unseen Environments
Figure 5 and Table 1 present our exploration experiment results. Across all settings, CARE enables robots to travel $2.9\times$ to $10.7\times$ longer before the first collision compared to NoMaD and ViNT without CARE. These performance variations correlated with the camera specifications of each configuration and the resulting depth estimation quality.
@@
LoCoBot with CARE exhibits both the longest distance traveled and the most significant improvement ($10.7\times$). This performance increase can be attributed to two main factors: the wide-angle fisheye camera enables extensive environmental perception, effectively detecting peripheral obstacles; and NoMaD, which has been predominantly trained on fisheye camera data, performs better with the camera setup of LoCoBot, which CARE further enhances.
TurtleBot4 with CARE exhibits moderate improvement ($3.3\times$) despite having the lowest distance traveled. This enhancement is partly due to its camera mounting position. When navigating corners, the rear-mounted camera maintains visibility of obstacles throughout the turn, allowing CARE to make safer adjustments. However, its narrow FOV still limits overall performance, particularly when approaching walls head-on.
RoboMaster without CARE achieves higher distance traveled due to its wider FOV, but shows the smallest improvement ($2.9\times$) with CARE. This limited gain stems from its front-mounted camera. During turning, the camera moves past obstacles before the body of the robot completes the turn, creating blind spots where side collisions can occur without detection. While the wider FOV achieves relatively higher distance traveled, it cannot fully compensate for these cornering blind spots.
@@ 4.2 Image Goal-Conditioned Navigation with Unseen Obstacles
We evaluate NoMaD and ViNT and their CARE-integrated variants in image goal-conditioned navigation scenarios involving unseen obstacles with the three robot platforms. We first establish a topological graph in a corridor environment to enable global planning. During testing, we place several OOD obstacles not represented in the topological map, creating scenarios that would typically require model fine-tuning to handle appropriately. Table 2 presents comprehensive results comparing the baseline navigation models (NoMaD and ViNT) against the same models integrated with CARE. All experiments are conducted across 10 distinct obstacle settings, with randomized positions for each trial. [^3] Detailed information about obstacle placements for each trial is available in Appendix B.1.
@@
Navigation with dynamic obstacles. We conducted additional tests in the same environment as the goal-conditioned experiments (see Figure 4(b)), where up to three people abruptly crossed the robot path during navigation. A trial was considered successful if the robot reached the goal without collision. Unlike prior works, we introduced dynamic obstacles that appeared suddenly from outside the camera field of view. NoMaD often failed to react, while CARE-integrated robots consistently exhibited reactive stopping and safe avoidance, demonstrating the ability to handle unpredictable dynamic obstacles where vision-only models without fine-tuning often fail to generalize. Further details are provided in Appendix A.1.
These results demonstrate the consistent ability of CARE to enhance navigation safety across various robot platforms without additional training. While performance varies based on camera configuration, CARE considerably improves collision avoidance in all tested scenarios, augmenting the capabilities of baseline models to handle complex environments with OOD objects. Supplementary videos of these experiments are available on our project page: https://airlab-sogang.github.io/CARE/
@@ 6 Limitations
7 Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00461409), Korea Planning & Evaluation Institute of Industrial Technology (KEIT) grant funded by the Korea government (MOTIE) (RS-2024-00444344), and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2022-00143911, AI Excellence Global Innovative Leader Education Program).
@@ B.3 Common Parameters for CARE
$\epsilon$ (m) 0.05-0.05 Vertical offset for filtering out ceiling points during top-down projection. A negative value indicates upward exclusion